Please help me to solve this: 3e^-x - 4 = 9
thanks
*Triston

Answer by Nate(3500)

3e^-x - 4 = 9 original
3e^-x = 13 add four
e^-x = (13/3) divide by 3
-x = ln (13/3) take natural log of each side
x = - ln (13/3) multiply by negative one
x is about -1.4663